OpenStudy (anonymous):

For f(x) = x^2 – 3x + 1, find all the values of c on the interval (0, 1) that satisfy the Mean Value Theorem. I keep getting 2???

OpenStudy (anonymous):

I do f(b) - f(a) / b - a so I get -1-1/1-0 so I get -2/1? so I get -2? I know this is wrong but its the answer I keep getting

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can not be done because the endpoints are not defined. Unless you meant [0,1]

OpenStudy (anonymous):

this is how it was written....

OpenStudy (loser66):

Assume that the interval [0,1], you have take int on that interval

OpenStudy (anonymous):

then i would say the hypotheses does not meet, therefor, can not be done

OpenStudy (anonymous):

int?

OpenStudy (jdoe0001):

the values of "x" between 0. .... 1 to see around what point it changes sign

OpenStudy (anonymous):

between .32 and .39

OpenStudy (jdoe0001):

using that table, yes

OpenStudy (anonymous):

okay sooo now what?

OpenStudy (jdoe0001):

now .... you found them =)

OpenStudy (jdoe0001):

the idea is to make a table of small values in between, to spot the sign change

OpenStudy (anonymous):

okay so my options are c = 1 c = - 0.5 c = 0.5 c = 0

OpenStudy (jdoe0001):

hmmm are you supposed to pick one?

OpenStudy (anonymous):

yeah... that's how I knew my -2 answer was wrong

OpenStudy (jdoe0001):

then it's no... that one,.... dunno

OpenStudy (ybarrap):

$$ f(x) = x^2 – 3x + 1\\ f'(x)=2x-3\\ $$ MVT, says that there exists a c such that $$ f'(c)=\cfrac{f(b)-f(a)}{b-a} $$ Let's find that c. The average slope between b and a is $$ \cfrac{-1-1}{1-0}=-2 $$ So then we need \(f'(c)=-2\): $$ f'(c)=2c-3=-2\\ \implies 2c=-2+3\\ \implies c=\cfrac{1}{2} $$ That it!